A 2500 kg delivery truck contains 125 cavans of rice, each of mass 500 kg. Its moving at a speed of 5m/s. How many cavans of rice must be unloaded at a nearby grocery so that the track can move at 9m/s and yet maintain the same momentum( Identify the given, unknown and equation to use.)

1
Expert's answer
2022-01-18T10:24:36-0500

Initial momentum:


"p_i=(M+N_1m)v_i,\\
p_f=(M+N_fm)v_f,\\
p_i=p_f,\\\space\\
(M+N_1m)v_i=(M+N_fm)v_f,\\\space\\
N_f=\dfrac{(M+N_1m)v_i-Mv_f}{mv_f}=71,\\\space\\
\Delta N=N_i-N_f=54."